### 03 Mar 2026 v1

# Implicit Bias in Deep Linear Discriminant Analysis

Abstract
While the Implicit Bias(or Implicit Regularization) of standard loss functions has been studied, the optimization geometry induced by discriminative metric-learning objectives remains largely unexplored.To the best of our knowledge, this paper presents an initial theoretical analysis of the implicit regularization induced by the Deep LDA—a scale-invariant objective designed to minimize intra-class variance and maximize inter-class distance. By analyzing the gradient flow of the loss on a L-layer diagonal linear network, we prove that under balanced initialization, the network architecture transforms standard additive gradient updates into multiplicative weight updates, which demonstrates an automatic conservation of the $||\cdot||_{2/L}$quasi-norm.
keywords:
Implicit Bias, LDA, Optimization, Deep learning
1 Introduction
Deep learning had already been successfully applied in numerous region. In the realm of deep learning, researchers widely believed there’s an unknown implicit factor that plays as a constraint in the training process of gradient-based models, ensuring their weights do not explode, which we called as “Implicit Bias”. With the involvement of more real-time applications, the subject of implicit bias become a more extensive direction in recent years. However, these theoretical breakthroughs are still majorly confined to loss functions with exponential tails (eg, Cross-Entropy) or square losses.
In classical statistical learning, Fisher Discriminant Analysis(FDA) is a classical method get proposed by R.A. Fisher in 1936, aiming for preserve information of data after reducing dimensions in a classification task [3]. In Matthias Dorfer’s group’s works around 2015 , they propose a new loss function based on FDA, Deep Linear Discriminant Analysis(Deep LDA) [2], with the same design concept, to maximize the inter-class distance while minimizing the intra one. While empirical studies demonstrate that such discriminative objectives yield highly separable features, their Implicit Bias still remains largely unexplored.
2 Literature Review
Implicit Bias, as previously mentioned, is one of the main hypothesis try to explain models’ good generalization performance, assuming there exists an implicit inclination of the model that select features. Understanding them better could lead more mathematical rigorous design to the loss function, and further imply the future improvement of model structures. As the left subplot show in figure.1, the gradient decent have an inclination that lead a same solution. Instead of using the traditional method to direct discuss its bound and convergence, the Implicit Bias provides another framework to analyze the reasons causing it in terms of the dynamics between features that tend to be bound by a constraint. Just as the right subplot illustrates, the deeply learned features repel each other and unfold from the origin,which makes a regular tetrahedron.
Figure 1: The Optimization in Simplex and the Implicit Bias
As for discussing the implicit regularization in logistic functions, the most famous research is Daniel Soudry’s work in 2018, finding that the weight vector is approaching infinity while still approximating the direction of $||\cdot||_{2}$ norm [8]. This result explains why the generalization performance still improves even after the loss function becomes stable in classification tasks. After a while, the conclusion from Ziwei Ji and Matus Telgarsky discover for logistic-type loss functions, the gradient descent would lead to self-alignment between the weights of each layer. The square loss is the most common type of loss function, hence it gets analyzed earlier. Suriya Gunasekar’s group focuses on matrix decomposition,and accidentally got the result that certain initialization could lead model weights to low-rank[6]. Just in next year, the same Suriya Gunasekar find the square loss in CNN with a stride would trigger sparsity in domain frequency[5].
Albeit the previous works talk about all prove to be have an Implicit Bias, but not loss function could prove to have an explicit mathematical Implicit Bias as Noam Razin suggested [7]. Building on Razin’s observation, it becomes imperative to explore whether a fundamentally different class of objectives could reveal entirely new optimization geometries. For instance, Deep LDA, as a famous loss adopting in wide range of applications especially in Compurter Vision(for example, Deep LDA use in COVID-19 diagnose[9]), but the Implicit Bias of it still remains as an open question. To fill the gap, our work formulates the gradient flow of the Rayleigh Quotient, revealing that this specific metric-learning objective explicitly imposes a strict quasi-norm conservation.
3 Methodology
We further extend this implicit bias analysis to a scale-invariant situation and try to figure out why and how Deep LDA induces feature sparsity. When knowing the specific Rayleigh Quotient objective, the gradient flow could be further simplified and reveal the whole optimization process by using continuous-time dynamics, in that way, if the reduction of redundant weights is happening with the more severe multiplicative penalty that network depth brought, exact quasi-norm conservation could get work.
Diagonal Linear Networks(DLNs) is a standard proxy model to discuss the model regularization [1], due to its special property allowing analysis focus on depth of network without other distributions. Hence for our theoretical setup, we adopt an $L$-layer DLN to rigorously analyze the gradient flow dynamics of the Deep LDA objective. In this paper, the research only discusses the continuous gradient flow under balanced initialization, not including non-linear activation functions or other complex network architectures.
3.1 Diagonal Linear Networks(DLNs)
First, we defined a deep linear network with $L$ layers representing the mapping from input to output. $u^{(1)}$ represent the weights of the first layer, and $u^{(L)}$ is the last one. In addition, all of them are restricted to diagonal matrices. Those sequences of weights are not fully connected in the definition, which means we are not just simply defining dense matrices but isolated parallel paths for each feature dimension.
$w_{i}=\prod_{k=1}^{L}u_{i}^{(k)}$ (1)
3.1.1 Conservation Lemma for DLNs
We need to prove there exist a conservation for DLNs first, for a better overview and to reduce the computation for Deep LDA in the next subsection. For any ${k}^{th}$ layer, we could get the effects each element in this layer if it gets impacted by the weight $w$ using the chain rule:
$\frac{\partial\mathcal{L}}{\partial u_{i}^{(k)}}=\frac{\partial\mathcal{L}}{\partial w_{i}}\frac{\partial w_{i}}{\partial u_{i}^{(k)}}=\frac{\partial\mathcal{L}}{\partial w_{i}}\left(\prod_{m\neq k}u_{i}^{(m)}\right)=\frac{\partial\mathcal{L}}{\partial w_{i}}\frac{w_{i}}{u_{i}^{(k)}}$ (2)
Assume variable $t$ is the current step, for standard gradient descent, the expression for the element $u_{k}$ would be like:
$u_{i}^{(k)}(t+1)=u_{i}^{(k)}(t)-\eta\frac{\partial\mathcal{L}}{\partial u_{i}^{(k)}}$ (3)
In eq.(3), we transform it into a form easy to become derivative about $t$. First move $u_{i}^{(k)}(t)$ to the left, and then divide $\eta$ in same time.
$\frac{u_{i}^{(k)}(t+1)+u_{i}^{(k)}(t)}{\eta}=-\frac{\partial\mathcal{L}}{\partial u_{i}^{(k)}}$ (4)
For getting the analytical solution, we take the limit of the eq.(4),and assume $\eta\rightarrow 0$. When $t$ also approaches 0, according to the definition of the derivative, the expression eq.(4) get convert to a gradient flow in eq.(5).
$\lim_{\eta\to 0}\frac{u_{i}^{(k)}(t+\eta)-u_{i}^{(k)}(t)}{\eta}=\frac{du_{i}^{(k)}}{dt}=-\frac{\partial\mathcal{L}}{\partial u_{i}^{(k)}}$ (5)
Substitute $\frac{\partial\mathcal{L}}{\partial u_{i}^{(k)}}$ in eq.(2) as $\frac{du_{i}^{(k)}}{dt}$, we could get eq.(6).
$\frac{du_{i}^{(k)}}{dt}=-\left(\frac{\partial\mathcal{L}}{\partial w_{i}}\frac{w_{i}}{u_{i}^{(k)}}\right)$ (6)
Next step, we multiply right hand side(RHS) and left hand side(LHS) in same time to move the denominator to the left:
$u_{i}^{(k)}\frac{du_{i}^{(k)}}{dt}=-w_{i}\frac{\partial\mathcal{L}}{\partial w_{i}}$ (7)
Due to the basic calculus rule $\frac{1}{2}\frac{d}{dt}(x^{2})=x\frac{dx}{dt}$, we could further simplify it as form in eq.(8).
$\frac{1}{2}\frac{d}{dt}\left(u_{i}^{(k)}\right)^{2}=-w_{i}\frac{\partial\mathcal{L}}{\partial w_{i}}$ (8)
By observing the eq.(8), we could see the whether what $k^{th}$ layer is, the right hand side is keep as constant. Hence we could get $\frac{d}{dt}\left(u_{i}^{(k)}\right)^{2}=\frac{d}{dt}\left(u_{i}^{(m)}\right)^{2}$.
And take the integral to both sides:
$\left(u_{i}^{(k)}(t)\right)^{2}-\left(u_{i}^{(m)}(t)\right)^{2}=C$ (9)
Furthermore, we utilize balanced initialization, assume $u_{i}^{(1)}(0)=u_{i}^{(2)}(0)=\dots=u_{i}^{(L)}(0)$,hence for $t=0$, there always be $\left(u_{i}^{(k)}(t)\right)^{2}=\left(u_{i}^{(m)}(t)\right)^{2}$. Additionally, the weight is always positive, so we could further conclude:
$u_{i}^{(1)}(t)=u_{i}^{(2)}(t)=\dots=u_{i}^{(L)}(t)$ (10)
As we defined in eq.(1), the form of $w_{i}(t)$ could write as :
$w_{i}(t)=\left(u_{i}^{(k)}(t)\right)^{L},\ u_{i}^{(k)}(t)=w_{i}(t)^{\frac{1}{L}}$ (11)
3.2 Implicit Bias in Deep LDA
$\mathcal{L}(w)=\frac{w^{\top}S_{w}w}{w^{\top}S_{b}w}$ (12)
The eq.(12) shows the main concept of this scale-invariant objective, it assumes the loss of the whole dataset could be formulated as formed in a Rayleigh Quotient. We express the loss as the ratio of the intra-class variance and the inter-class variance, as Deep LDA(or Generalized Rayleigh Quotient) is defined, when $w\neq 0$ and $w^{\top}Bw\neq 0$.
When we multiply the $w$ by the factor $a$, the loss function of Deep LDA would be:
$\mathcal{L}(\alpha w)=\frac{\alpha w^{\top}A(\alpha w)}{\alpha w^{\top}B(\alpha w)}=\frac{\alpha^{2}(w^{\top}Aw)}{\alpha^{2}(w^{\top}Bw)}$ (13)
The $a^{2}$ gets offset, we could get $\mathcal{L}(\alpha w)=\frac{w^{\top}Aw}{w^{\top}Bw}=\mathcal{L}(w)$, which denotes that Deep LDA is a Homogeneous function of degree 0, this is the first finding.
Tracing back the theoretical setup we did in section 3.1.1. We prove $\frac{\partial w_{i}}{\partial u_{i}^{(k)}}=\frac{w_{i}}{u_{i}^{(k)}}$(simply take partial derivative) and $\frac{du_{i}^{(k)}}{dt}=-\frac{\partial\mathcal{L}}{\partial w_{i}}\left(\frac{w_{i}}{u_{i}^{(k)}}\right)$. So the total derivative $\frac{dw_{i}}{dt}$ could convert to an expression about the loss function in eq.(14).
$\frac{dw_{i}}{dt}=\sum_{k=1}^{L}\frac{\partial w_{i}}{\partial u_{i}^{(k)}}\frac{du_{i}^{(k)}}{dt}=\sum_{k=1}^{L}\left(\frac{w_{i}}{u_{i}^{(k)}}\right)\left[-\frac{\partial\mathcal{L}}{\partial w_{i}}\left(\frac{w_{i}}{u_{i}^{(k)}}\right)\right]=-\frac{\partial\mathcal{L}}{\partial w_{i}}\sum_{k=1}^{L}\frac{w_{i}^{2}}{\left(u_{i}^{(k)}\right)^{2}}$ (14)
Because $u_{i}^{(k)}(t)$ get the result is $w_{i}(t)^{\frac{1}{L}}$ under balance initialization,we could replace the denominator.
$\frac{dw_{i}}{dt}=-\frac{\partial\mathcal{L}}{\partial w_{i}}\sum_{k=1}^{L}\frac{w_{i}^{2}}{w_{i}^{\frac{2}{L}}}=-L\cdot w_{i}^{2-\frac{2}{L}}\frac{\partial\mathcal{L}}{\partial w_{i}}$ (15)
Since we prove Scale Invariance in eq.(13),and its have natural property in eq.(16).(see Appendix A or the book Matrix Computations[4] for the complete proof, in geometrical aspect to understand, only when it orthogonal to the direction of weight could cause it not effected by scaling)
$w^{\top}\nabla_{w}\mathcal{L}=\sum w_{i}\frac{\partial\mathcal{L}}{\partial w_{i}}=0$ (16)
And we further multiply the both sides of the eq.(15) with $w_{i}^{\frac{2}{L}-1}$.
$w_{i}^{\frac{2}{L}-1}\frac{dw_{i}}{dt}=-L\cdot w_{i}^{\left(2-\frac{2}{L}\right)+\left(\frac{2}{L}-1\right)}\frac{\partial\mathcal{L}}{\partial w_{i}}w_{i}^{\frac{2}{L}-1}=-L\cdot w_{i}\frac{\partial\mathcal{L}}{\partial w_{i}}$ (17)
Taking the integral again to both sides of eq.(17):
$\frac{d}{dt}\left(w_{i}^{\frac{2}{L}}\right)=-2w_{i}\frac{\partial\mathcal{L}}{\partial w_{i}}$ (18)
$\frac{d}{dt}\sum_{i=1}^{d}w_{i}^{\frac{2}{L}}=-2\sum_{i=1}^{d}w_{i}\frac{\partial\mathcal{L}}{\partial w_{i}}=-2w^{\top}\nabla_{w}\mathcal{L}(w)$ (19)
Because Scale Invariance in eq.(16),we could get an expression for the sum up of weights in all dimensions is get blanced and bound with a constant $C$.
$\sum_{i=1}^{d}w_{i}(t)^{\frac{2}{L}}=C$ (20)
To other words, the whole training process, the Deep LDA is bounded by the |w(t)|2/L|w(t)|{2/L} as the expression show beneath:
$|w(t)|{2/L}^{2/L}=|w(0)|_{2/L}^{2/L}\quad\forall t\geq 0$ (21)
4 Experiment
Our experiment settled a test in DLNs implemented with the basic Numpy package, testing with hyper-parameters( $\eta=0.005$ and $100000$ epochs). The test is mainly for comparing the dynamic differences between the networks with different numbers of layers $L$, including $L=[1,2,5,10,20]$. For the experimental setup, we defined a simplified feature space with a dimension of $d=5$, and synthesized the intra-class and inter-class scatter matrices ($S_{w}$ and $S_{b}$) as random positive-definite matrices from a uniform distribution $\mathcal{U}(0.4,0.6)$.
In the top-left subplot of fig.2, we find five horizontal plain lines,which validate our conclusion in eq.(9), no matter how many layers a DLN gets, the implicit regularization remains unchanged.
About the second subplot to the last one, the ones about weight evolution, it clearly denotes with an increasing number of layers. With a smaller number of layers, the weak features get eliminated faster than the DLNs, and slower for the strong features converge. The weak features are selected out faster under the expectation due to our conclusion in (21). The phenomenon that strong features get fluctuating is due to the Edge of Stability caused by our stable learning rate,since in previous conduct we were all using $dt$ rather than a constant learning rate.
Figure 2: The Simulated Result for DeepLDA in DLNs.
5 Conclusion
In conclusion, we provide a theoretical perspective on the implicit bias induced by Deep Linear Discriminant Analysis (Deep LDA) under a diagonal linear network (DLN) framework. Crucially, this conservation stem from two key ingredients: (1) the multiplicative parameterization induced by depth, and (2) the scale-invariance property of the Rayleigh quotient objective. Together, these yield a strict geometric constraint on the optimization trajectory. The experiment finds that deeper architectures amplify multiplicative penalties on weak features, promoting sparsity-like behavior in the effective weights.
Our current findings are fundamentally constrained by the simplified diagonal linear architecture without non-linear activations, need further validations for real-world data. Future works will extend this continuous-time framework beyond diagonal networks to incorporate non-linear activation functions. Additionally, investigating how stochastic gradient descent impacts this strict quasi-norm conservation in Deep LDA remains a critical open direction.
Appendix A Proof of the Orthogonality Lemma
In this section, we provide the rigorous mathematical derivation for the gradient of the Deep LDA objective and prove the orthogonality property $w^{\top}\nabla_{w}\mathcal{L}(w)=0$ referenced in Section 3.2.
Let the generalized Rayleigh Quotient be defined as:
$\mathcal{L}(w)=\frac{N(w)}{D(w)}=\frac{w^{\top}S_{w}w}{w^{\top}S_{b}w}$ (22)
where the numerator is $N(w)=w^{\top}S_{w}w$ and the denominator is $D(w)=w^{\top}S_{b}w$.
A.1 Gradient of the Rayleigh Quotient
Because both the intra-class scatter matrix $S_{w}$ and the inter-class scatter matrix $S_{b}$ are symmetric real matrices, the derivative of their associated quadratic forms with respect to the vector $w$ can be straightforwardly evaluated as:
$\nabla_{w}N(w)=2S_{w}w,\quad\nabla_{w}D(w)=2S_{b}w$ (23)
Applying the standard quotient rule for differentiation, $\nabla\left(\frac{N}{D}\right)=\frac{D\nabla N-N\nabla D}{D^{2}}$, we obtain the gradient of $\mathcal{L}(w)$:
$\nabla_{w}\mathcal{L}(w)=\frac{(w^{\top}S_{b}w)(2S_{w}w)-(w^{\top}S_{w}w)(2S_{b}w)}{(w^{\top}S_{b}w)^{2}}$ (24)
By factoring out $\frac{2}{w^{\top}S_{b}w}$, we can rewrite the gradient elegantly in terms of $\mathcal{L}(w)$ itself:
$\nabla_{w}\mathcal{L}(w)=\frac{2}{w^{\top}S_{b}w}\left(S_{w}w-\left(\frac{w^{\top}S_{w}w}{w^{\top}S_{b}w}\right)S_{b}w\right)=\frac{2}{w^{\top}S_{b}w}\big(S_{w}w-\mathcal{L}(w)S_{b}w\big)$ (25)
A.2 Scale Invariance Property
Using the derived gradient in Eq. (25), we can now calculate the inner product between the weight vector $w$ and its gradient $\nabla_{w}\mathcal{L}(w)$:
$w^{\top}\nabla_{w}\mathcal{L}(w)=w^{\top}\left[\frac{2}{w^{\top}S_{b}w}\big(S_{w}w-\mathcal{L}(w)S_{b}w\big)\right]$ (26)
Since $\frac{2}{w^{\top}S_{b}w}$ is a scalar, we distribute $w^{\top}$ into the parenthesis:
$w^{\top}\nabla_{w}\mathcal{L}(w)=\frac{2}{w^{\top}S_{b}w}\Big(w^{\top}S_{w}w-\mathcal{L}(w)w^{\top}S_{b}w\Big)$ (27)
Substituting the original definition of $\mathcal{L}(w)=\frac{w^{\top}S_{w}w}{w^{\top}S_{b}w}$ back into the second term:
$w^{\top}\nabla_{w}\mathcal{L}(w)=\frac{2}{w^{\top}S_{b}w}\left(w^{\top}S_{w}w-\left(\frac{w^{\top}S_{w}w}{w^{\top}S_{b}w}\right)w^{\top}S_{b}w\right)$ (28)
The scalar $w^{\top}S_{b}w$ cancels out perfectly, leaving:
$w^{\top}\nabla_{w}\mathcal{L}(w)=\frac{2}{w^{\top}S_{b}w}\Big(w^{\top}S_{w}w-w^{\top}S_{w}w\Big)=0$ (29)
This completes the proof that the gradient of the scale-invariant Deep LDA objective is always orthogonal to the weight vector itself.
Competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding Information
Not Applicable
Author Contribution
Not Applicable
Data Availability Statement
The data were generated using np.random.randn with a fixed seed (8086) for reproducibility. Upon publication, the experiment will be available at https://www.kaggle.com/code/spike8086/implicit-bias-in-deep-lda.
Research Involving Human or Animals
Not Applicable
Informed Consent
Not applicable.